OpenStudy (anonymous):

What is the solution to (4x+8)=-8x+3

OpenStudy (anonymous):

Are you looking for x?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Okay, so you want to start out with both variables on the same side. (4x+8)=-8x+3 4x + 8x + 8 = -8x + 8x + 3 12x + 8 = 3 12x = 5 x = 5/12 or 0.4166

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Sure thing.